OpenStudy (anonymous):

an airplane that travels at 550 mph in still air encounters a 50 mph headwind. how long will it take the plane to travel 900 mi in to the wind

OpenStudy (anonymous):

How long before it encounters the headwind? Is it asking how long it takes to travel 900 mi? or just in the time it encounters the headwind?

OpenStudy (jdoe0001):

sounds like against the headwind

OpenStudy (jdoe0001):

anyhow, distance = rate * t, against the wind the rate should be 550-50 distance of 900 solve for "t"

OpenStudy (anonymous):

So would i take 500 * T

OpenStudy (anonymous):

distance=rate*t as said above.. 900=500*t divide 900 by 500, and you get 1.8. 1.8=108 minutes since its mph.

OpenStudy (anonymous):

1= 60 min you can calculate what 0.8 is and add the two together. :)

OpenStudy (anonymous):

Omg thank you both I am helping my sister with her college school work